OpenStudy (anonymous):

write the following equations without logarithms a) logT= 2logx-log y b) log(base2) K= log(base2)n +1/2log(base2)t

OpenStudy (anonymous):

a) logT = logx^2/y T = x^2/y b) log(base2)K = log(base2)nt^(1/2) K = nt^(1/2)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw